You are planning to acquire a business today and sell it in 3 years. The cash flows associated with this investment are: 219,496 an expense of dollars in year zero; an expense of 36,201 dollars in year 1; a receivable of 27,681 dollars in year 2; a receivable of 38,789 dollars in year 3. In addition, you expect to sell the business for 364,064 dollars upon the last receivable. If the MARR (Minimum Acceptable Rate of Return) is 5% per year, compute the annual worth of this investment.
Answer #1

ANSWER:

I = 5%

N = 3

PW = Cash flow in year 0 + cash flow in year 1(p/f,i,n) + cash flow in year 2(p/f,i,n) + cash flow in year 3(p/f,i,n)

PW = -219,496 - 36,201(p/f,5%,1) + 27,681(p/f,5%,2) + 38,789(p/f,5%,3)

PW = -219,496 - 36,201 * 0.9524 + 27,681 * 0.907 + 38,789 * 0.8638

PW = -219,496 - 34,477.8324 + 25,106.667 + 33,505.9382

PW = -195,361.2272

AW = PW(A/P,I,N)

AW = -195,361.2272(A/P,5%,3)

AW = -195,361.2272 * 0.3672

AW = -71,736.64

SO THE ANNUAL WORTH OF THIS INVESTMENT IS -$71,736.64